### 12 Mar 2026 v1

# Infrared Thermography in the Tokamak à Configuration Variable

Abstract
In the Tokamak à Configuration Variable (TCV), infrared thermography (IR) is currently composed of the horizontal, vertical, and tangential infrared systems (HIR, VIR, TIR), which all use Equus 81k M cameras. The IR diagnostics obtain the surface temperature of TCV’s graphite tiles for post-discharge analysis. Target heat flux profiles are inferred from the tile temperature with the THEODOR (Thermal Energy Onto Divertor) code. Fast transient analysis is possible in reduced frame mode, with acquisition frequencies above 10kHz. The main views are the lower inner wall for HIR, the floor for VIR, and the lower outer wall for TIR. The HIR camera can also be moved to view the midplane inner wall, while TIR can be moved to see the midplane inner wall and the upper outer wall, mainly to measure synchrotron radiation and heat deposition due to runaway electrons. Recent developments in TCV’s IR systems include (i) tile diffusivity and conductivity measurements to assure the precision of heat flux estimates; (ii) the addition of one new VIR heated valley tile and two rooftop TIR tiles, for measurements of fast heat flux transients; (iii) the implementation of long-pass wavelength filter of 4095 nm, to diminish the measurement of plasma parasitic infrared light, mainly from deuterium $5\rightarrow 4$ emission at 4051 nm. Despite these developments, the main sources of uncertainty for IR in TCV are still parasitic infrared light and the determination of the surface layer heat transmission factor, both of which mainly affect the VIR system.
École Polytechnique Fédérale de Lausanne (EPFL), Swiss Plasma Center (SPC),
CH-1015 Lausanne, Switzerland
1 Introduction
Infrared thermography (IR) is an imaging technique that translates the infrared radiation emitted by an object into its surface temperature. This is done by employing the total photon radiance from blackbody theory, $R_{BB}$ [1]:
$R_{BB}(T)=\int_{\lambda_{1}}^{\lambda_{2}}\frac{2c}{\lambda^{4}}\frac{d\lambda}{\exp\left(\frac{hc}{\lambda k_{B}T}\right)-1}\quad\left[\frac{\text{1}}{\text{sr}\;\text{s}\;\text{m}^{2}}\right]$ (1.1)
where $c=2.99792\times 10^{8}\;\text{m/s}$ is the light speed, $h=6.62607\times 10^{-34}\;\text{J s}$ is Planck’s constant, $k_{B}=1.38065\times 10^{-23}\;\text{J/K}$ is Boltzmann’s constant, $T$ is the object surface temperature, and $\lambda_{1}$ and $\lambda_{2}$ are the minimum and maximum wavelength measured by the camera detector. $R_{BB}(T)$ measures the number of photons per solid angle, time and area emitted by the object. Inverting $R_{BB}(T)$ gives the surface temperature as a function of the infrared radiation. In tokamaks and stellarators, infrared thermography is used to evaluate the temperature of the device walls in contact with the plasma. By employing heat diffusion equations, the heat flux due to the plasma can then be estimated.
This paper details the IR systems of the Tokamak à Configuration Variable (TCV) [2]. Sec. 2 describes TCV’s infrared cameras; Sec. 3, the calibration to temperature with thermocouples; Sec. 4, the target heat flux estimation with the THEODOR code; Sec. 5, the windows used for IR in TCV; Sec. 6, the tiles — with emphasis for their thermal properties and new tiles developed for fast transient analysis; Sec. 7 discusses the infrared parasitic plasma radiation measured by the cameras; Sec. 8 shows a comparison between the VIR and TIR heat flux profiles and Sec. 9 summarizes the results.
2 TCV’s infrared cameras
2.1 General information
All infrared cameras in TCV are currently IRCAM Equus 81k M (specifications in Table 2.1), adapted by the manufacturer to withstand ‘high’ magnetic fields (0.1-0.2 T). The detector is a focal plane array (FPA) made of mercury cadmium telluride (MCT). A Stirling cooler refrigerates the detector, minimizing its own infrared emission and reducing noise due to thermally excited electron holes. The sensor is a complementary metal oxide semiconductor (CMOS), with each pixel having its own amplifier. The length of each pixel (i.e. the pixel pitch) is of $30\;\mu\text{m}$. The integration mode of the data is integrate-then-read.
The cameras’ spectral response goes from $3.7\;\mu\text{m}$ to $4.8\;\mu\text{m}$, falling in the mid-wavelength infrared (MWIR) range. Nevertheless, as of TCV’s discharge 79993, all three cameras are filtered with a $4095\;\text{nm}$ long-pass wavelength filter to decrease the amount of infrared light measured from the plasma.
The maximum and minimum frame sizes are respectively 320x256 (approximately 81 thousand pixels, hence the camera name Equus 81k M) and 64x2 pixels. Decreasing the frame size increases the acquisition frequency, as described in Eq. (2.1). The spatial resolution ranges from about 2.0mm/pixel (VIR and TIR) to 1.3mm/pixel (HIR).
The minimum integration time (i.e. the interval the camera pixels are exposed to light) is $\Delta t_{int}=1\;\text{$\mu$s}$, which can be used for cases with very high tile temperatures (e.g. HIR of discharge 89000, not shown). Nevertheless, the standard value is $\Delta t_{int}=1\;\text{ms}$ for L-mode, ohmically heated plasmas.
TCV’s infrared cameras are controlled by a C++ program with MDS events (Sec. 3.1 of [3]). As of 2025, the quality of the data is monitored shot by shot in TCV’s DdJ panel [4].
Table 2.1: General information about TCV’s IR cameras Equus 81k M by IRCAM.
Detector MCT FPA
Sensor CMOS
Integration mode Integrate-then-read
Wavelength sensitivity $3.7\;\mu\text{m}$ (or $4.1\;\mu\text{m}$ filtered) to $4.8\;\mu\text{m}$
Max window size 320x256 pixels
Min window size 64x2 pixels
Pixel pitch $30\;\mu\text{m}$
Spatial resolution 1.3mm/pixel (HIR) and 2mm/pixel (VIR & TIR)
Integration time $1\;\mu\text{s}$ to $10\;\text{ms}$
Acquisition frequency 200 Hz (standard) to 20kHz (max)
2.2 Interval between frames
As described in the cameras user manual and in Sec. 3.1.2 of [3], the minimum interval between frames of TCV’s 81k M IR cameras is
$\Delta t_{frame}^{(min)}$ $=\Delta t_{int}+\Delta t_{read}+\Delta t_{reset}$ (2.1)
$=\Delta t_{int}+N_{px}/(4f_{clk})+20/f_{clk}$
where $\Delta t_{int}$ is the integration time, i.e. the interval in which the camera receives light for each frame. $\Delta t_{read}=N_{px}/(4_{clk})=N_{px}/27$ is the readout time of the data, $N_{px}$ is the number of pixels used by the camera (max $320\times 260=81920$, min $64\times 2=128$, as in Table 2.1) and $f_{clk}=6.75\;\text{MHz}$ is the camera mother-clock frequency. $\Delta t_{reset}=20/f_{clk}=20/(6.75\;\text{MHz})=3\;\mu\text{s}$ is estimated as the time to reset the pixels for the next measurements. It is also possible to change $f_{clk}$ to $10\;\text{MHz}$, decreasing $\Delta t_{read}$ by 32%. However, this leads to pixels interfering to each other (pixel cross-talk), which can distort the data.
Substituting $f_{clk}=6.75\;\text{MHz}$ in Eq. (2.1) gives the general formula for the minimum interval between frames
$\Delta t_{frame}^{(min)}=\Delta t_{int}+\frac{1}{27}N_{px}\;\mu\text{s}+3\;\mu\text{s}$ (2.2)
In full frame mode, $N_{px}=320\times 260=81920$, yielding $\Delta t_{read}=3.0\;\text{ms}$ and hence
$\Delta t_{full\;frame}^{(min)}=\Delta t_{int}+3.0\;\text{ms}$ (2.3)
Since in general $\Delta t_{int}<3\;\text{ms}$, the full frame interval is dominated by the readout time and the maximum acquisition frequency is then obtained for $\Delta t_{int}\rightarrow 0$, $f_{acq}=1/\Delta t_{full\;frame}^{(min)}=333\;\text{Hz}$.
In practice, to assure reliable measurements, the used interval between frames is increased by a factor of 10% and an adjustment $\Delta t_{adjust}$ is added,
$\Delta t_{frame}$ $=(\Delta t_{int}+\Delta t_{read}+\Delta t_{reset}+\Delta t_{adjust})\cdot 1.1$ (2.4)
$=(\Delta t_{int}+N_{px}/27\;\mu\text{m}+3\;\mu\text{s}+15\;\mu\text{s})\cdot 1.1$
$\Delta t_{adjust}=15\;\mu\text{s}$ is important for low number of pixels ($N_{px}<2400$), specially at extremely low integration times ($1\sim 5\;\mu\text{s}$) to assure the identification of the frame instant. Decreasing the readout time too much can also lead to frames not being recorded and thus a minimum of $N_{px}=64\times 24=1536$ is recommended.
2.3 HIR: the horizontal infrared system
The HIR standard view is the lower port of TCV’s graphite tiles, where the inner strike point (SP) of most of TCV’s plasmas is located (Fig. 2.2). HIR is also routinely moved to the midplane (HIRm), to measure the inner SP of long-legged configurations, synchrotron radiation from runaway electrons, and RE heat deposition. The history of HIR is described in Table 2.3.
Table 2.2: HIR history.
Date Event
15/7-7/9/09 HIR camera borrowed from MAST
2012 Decoupled HIR support from tokamak to avoid vibration
2013 Mid-wavelength camera used for HIR (1st shot: 48508)
2015 IRCAM Equus1 81k M camera acquired and now uses a new and a 25mm lens; re-allocated from TCV sector 4 to 7; can be placed on the midplane port or a lower port
Dec. 2015 Purchase of a 12.5mm lens from IRCAM
Jun. 2022 Midplane sapphire window replaced by germanium
23 Nov. 2022 Midplane germanium window cracks
Jan. 2023 New midplane sapphire window installed
Jul. 2025 6mm ZnSl lower port window with AR coating cracked on installation
25 Sep. 2025 New lower port sapphire window installed (as of discharge 87965)
Table 2.3: VIR history. 32951 is not the 1st discharge, but a discharge of interest.
Date Event 1st discharge
Oct. 2006 Thermosensorik camera installed 32951*
Mar. 2008 Optics v2, with better image quality and wider view 34756
Mar. 2017 IRCAM Equus2 81k M replaced Thermosensorik
09/10/20 VIR shutter broke
28/02/21 Ge window with AR coating replaced old ZnSe window 70129
Feb. 2023 Ge window replaced by sapphire after factor 3 decrease in transmissivity discovered 77871
09/11/23 M. Pedrini’s tile AG6831 replaces tiles 404, 405, and 406 78687
05/03/24 Equipped with a 4095 nm long wave pass filter 79993
23/07/25 Valley tile AH5251 replaces M. Pedrini’s tile AG6831 87125
Table 2.4: TIR history.
Date Event 1st discharge
06/06/19 1st discharge with TIR 63060
13/08/19 Port protection tiles installed in TCV 63528
12/07/22 Two long 45º tiles, AF7102 and AF7103, installed 74982
05/03/24 Equipped with a 4095 nm long wave pass filter 79993
20/03/25 Upper port support installed, for RE studies 85460
29/04/25 Lower port camera support rotated by 28º for ELM studies 85885
23/07/25 Rooftop tiles AH6513 & AH6514 (for ELM studies) replace tiles AF7102 & AF7103 87125
Figure 2.1: (a) Magnetic reconstruction of a power exhaust reference shot (87222), repeated from time to time to assess the divertor conditions, with all IR camera views except the TIR upper port. (b) HIR view, with inner strike point at the inner wall. Differences in temperature in the strike point are due to the inner wall tiles not being toroidally symmetric [5]. (c) VIR view with outer strike point (OSP) at the tilted tiles and line of interest (in red) used for the heat flux analysis in Fig. 8.1. Red and yellow circles are screws. Difference in temperature in the OSP are due to leading edges and absence of toroidal symmetry.
Figure 2.2: TIR main view before (a) and after (b) being rotated by 28º, during two identically programmed PEX reference shots with OSP at the tilted tiles. The dashed line indicates the line of interest used to estimate the target heat flux in Fig. 8.1. The discharge 87222 also already had the new rooftop tile installed. Higher temperature at the edges happens due to leading edges and the fact that the tilted tiles are not toroidally symmetric.
2.4 VIR: the vertical infrared system
VIR is the IR system most used in TCV, as it images the floor of TCV, where the outer strike point of TCV’s single null plasmas is usually localized. It also views the tilted tiles (Fig. 2.2). Originally, the VIR system had a Thermosensorik camera (for specification see Sec. 2.2.1.1 of [3]), which was replaced for an IRCAM Equus2 81k M in March 2017. VIR uses a relay optics with seven lenses (six of silicon, one of germanium) and a mirror (Sec. 4.3.1 of [6]). Table 2.3 describes a summary of VIR’s history.
2.5 TIR: the tangential infrared system
TIR is the newest thermography system in TCV (1st discharge: 63060, in 2019). Its main view maps the lower tilted tiles and port protection tiles (Fig. 2.2), which can be used to measure the outer strike point of alternative divertor configurations such as the super-X, X-point target, and negative triangularity plasmas. In December 2024, a new upper port view was installed for the TIR, mainly to measure synchrotron radiation and heat deposition due to REs. The camera can be rotated into 5 different positions in this view. The TIR history is summarized in Table 2.4. As of discharge 85885, the camera view has been rotated by 28º, to align its vertical direction with the tilted tiles, allowing smaller frame size and increased acquisition frequency.
3 Calibration to temperature
The model used in TCV to relate the IR camera digitized counts $N_{ij}$ at a given pixel $ij$ ($i\in[1,256]$ and $j\in[1,320]$) to the tile temperature $T_{ij}$ is
$N_{ij}(T_{ij},\Delta t_{int})=\left[a_{ij}\mathcal{I}(T_{ij})+b_{ij}\right]\Delta t_{int}+c_{ij}$ (3.1)
where $\mathcal{I}(T)\equiv R_{BB}(T)/(2c)$ (with $c$ being the light speed and $R_{BB}$ the total photon radiance, given by Eq. (1.1)), $a_{ij}\;[\text{counts sr m}^{3}/\text{ms}]$ is the effective transmittance of the signal, $b_{ij}\;[\text{counts/ms}]$ is a term accounting for dark currents and infrared emission from objects other than the tiles (e.g. lenses, windows), and $c_{ij}\;[\text{counts}]$ corresponds to the base level of the pixels. More details on the different effects that compose $a_{ij}$ can be found in Sec. 4.2 of [6].
To obtain the parameters $a_{ij}$, $b_{ij}$, $c_{ij}$, a two-step procedure is employed.
3.1 Step 1 — Heated tile
A heated tile embedded with a thermocouple (TC) is used to calibrate the camera signal to the tile temperature (Fig. 3.1),
$N_{TC}(T_{TC},\Delta t_{int})=A_{TC}(\Delta t_{int})\mathcal{I}(T_{TC})+B_{TC}(\Delta t_{int})$ (3.2)
where $N_{TC}$ is the average of the camera signal $N_{ij}$ over the closest pixels to the TC $(N_{TC}\equiv\left\langle N_{ij}\right\rangle_{TC})$ and
$A_{TC}(\Delta t_{int})$ $\equiv a_{TC}\Delta t_{int}$ (3.3)
$B_{TC}(\Delta t_{int})$ $\equiv b_{TC}\Delta t_{int}+c_{TC}$
with $a_{TC}\equiv\left\langle a_{ij}\right\rangle_{TC},$ $b_{TC}\equiv\left\langle b_{ij}\right\rangle_{TC},$ and $c_{TC}\equiv\left\langle c_{ij}\right\rangle_{TC}.$ Graphs of $N_{TC}(T_{TC},\Delta t_{int})$ vs $T_{TC}$ are obtained for several integration times (Fig. 3.2). Fitting them yields $A_{TC}(\Delta t_{int})$ and $B_{TC}(\Delta t_{int})$, which are then used to evaluate $a_{TC}$, $b_{TC}$, and $c_{TC}$ from fits of Eq. (3.3).
Figure 3.1: HIR raw data view during the temperature calibration with TC at $T=89\;\text{ºC}$. Heated tile (HT) in the center, with the pixels $ij$ used for the calibration in red.
Figure 3.2: (a) Raw HIR data vs TC temperature for $\Delta t_{int}=0.1\;\text{ms}$. (b) Fits of $A_{TC}$ and $B_{TC}$ vs $\Delta t_{int}$. Fitting Eq. (3.2) in Fig. 3.2(a) gives $A_{TC}$ and $B_{TC}$ at $\Delta t_{int}=0.1\;\text{ms}$. Doing the same for different integration times yields the data of Fig. 3.2(b), which can then be fitted by Eq. (3.3) to obtain $a_{TC}$, $b_{TC}$, and $c_{TC}$ (known as 3 parameter fit).
3.2 Step 2 — Non-uniformity correction
Step 1 obtains the calibration parameters at the thermocouple position. To evaluate $a_{ij},$ $b_{ij},$ and $c_{ij}$ for all other pixels, a non-uniformity correction is needed. For this, IR images of a blackbody (e.g. a piece of paper in front of the camera) at uniform temperature $T_{0}$ are taken at different integration times. The signal $N_{ij}(T_{0},\Delta t_{int})$ at each pixel can then be related to
$N_{ij}(T_{0},\Delta t_{int})=a^{\prime}_{ij}(T_{0})\Delta t_{int}+c_{ij}$ (3.4)
where $a^{\prime}_{ij}(T_{0})\equiv a_{ij}\mathcal{I}(T_{0})+b_{ij}$ (from Eq. (3.1)). Eq. (3.4) is then fitted at different integration times, yielding $a^{\prime}_{ij}$ and $c_{ij}$ for all pixels $i\in[1,256]$ and $j\in[1,320]$. The gain and offset indices for each pixel are then evaluated,
$g_{ij}$ $=\frac{a^{\prime}_{ij}}{a^{\prime}_{TC}}$
$o_{ij}$ $=\frac{c_{ij}}{c_{TC}}$
where $a^{\prime}_{TC}$ is the averaged $a^{\prime}_{ij}$ over the pixels around the thermocouple used in step 1. This results in the coefficients
$a_{ij}$ $=g_{ij}a_{TC}$
$b_{ij}$ $=g_{ij}b_{TC}$
$c_{ij}$ $=o_{ij}c_{TC}$
which are then used in Eq. (3.1) for each IR camera during TCV shots to translate the raw signal to surface temperature. The matrices of $a_{ij}$, $b_{ij}$, and $c_{ij}$ are shown in Fig. .
Figure 3.3: $a_{ij}$, $b_{ij}$, and $c_{ij}$ matrices for the HIR calibration shown in Fig. 3.2.
4 Heat flux estimation
4.1 From temperature to heat flux with THEODOR
The heat flux impinging perpendicularly to the tile is evaluated through Fourier’s law,
$\vec{q}_{\perp}=-k\nabla T$ (4.1)
where $k(T)$ is the tile conductivity and $T(x,y,t)$ is the tile temperature along a line of interest, with $x$ being the coordinate along the heat flux profile and $y$ the coordinate that goes from the surface to inside the tile. To obtain the temperature distribution in the tile, the heat diffusion equation is used,
$\rho c_{p}\frac{\partial T}{\partial t}=\nabla\cdot(k\nabla T)$ (4.2)
where $\rho=cte$ is the tile mass density and $c_{p}(T)$ is the tile specific heat. $\rho$, $c_{p}$, and $k$ are known. As it will be shown in Sec. 6.1, the $c_{p}$, and $k$ depend on the tile temperature, which complicates the solution of Eq. (4.2). However, introducing the heat flux potential
$U(T)\equiv\int_{0}^{T}k(T)dT,$ (4.3)
Eq. (4.2) can be rewritten as (see Appendix B of [7])
$\frac{\partial U}{\partial t}=D\nabla^{2}U$ (4.4)
where the tile heat diffusivity is defined as
$D(T)\equiv\frac{k(T)}{\rho c_{p}(T)}$ (4.5)
Eq. (4.4) is solved by the code THEODOR (Thermal Energy Onto Divertor) [8] to obtain $U$ as a function of time.
The following parametrization is used for the conductivity,
$k(T)=a_{k}+b_{k}\left(1+\frac{T}{T_{0,k}}\right)^{-2}$ (4.6)
where $a_{k}$, $b_{k}$ and $T_{0,k}$ are parameters to be determined with the available data. In TCV, $a_{k}=0$ was obtained (Sec. 6.1). The same functional dependence is used for the tile diffusivity,
$D(T)=a_{D}+b_{D}\left(1+\frac{T}{T_{0,D}}\right)^{-2}$ (4.7)
Integrating Eq. (4.3) using $k(T)$ from Eq. (4.6) yields
$U(T)=a_{k}T+b_{k}\frac{T}{1+T/T_{0,k}}$ (4.8)
Inverting this equation results in the tile temperature distribution, which is then used to obtain the heat flux perpendicular to the tile, $\vec{q}_{\perp}=-k\nabla T$ (Eq. (4.1)).
4.2 Boundary conditions and surface layer heat transmission factor
For IR in TCV, a python version of the THEODOR code is used, which solves the differential equation (4.4) implicitly. The three boundary conditions for this THEODOR version are
$\begin{cases}\left.\frac{\partial T}{\partial x}\right|_{\text{side}}=0\\ q_{\perp}(x,d_{tile},t)=0\\ T(x,0,t)=T_{layer}(x,t)-\frac{q_{\perp}(x,0,t-\Delta t)}{\alpha_{top}}\end{cases}$ (4.9)
The first equation means that THEODOR assumes that the temperature derivative at the edges of the heat flux profile is zero. The 2nd equation assumes that the heat flux at the bottom of the tile is zero.
Tile erosion and dust deposition can form surface layers that cause an overestimation of the actual tile temperature [8, 9, 10]. To correct for this effect and avoid false negative heat fluxes, a surface layer heat transmission factor is introduced,
$\alpha_{top}=\frac{k_{layer}}{d_{layer}}$ (4.10)
where $k_{layer}$ and $d_{layer}$ are respectively the conductivity and thickness of the layer. The third boundary condition relates the temperature of the surface layer $T_{layer}(x,t)$ (i.e., the temperature actually measured by the IR cameras), to the tile temperature just below the surface layer, $T(x,0,t)$. $q_{\perp}(x,0,t-\Delta t)$ is used since $q_{\perp}(x,0,t)$ is not known at the time where the boundary condition is computed.
Values for $\alpha_{top}$ in TCV range from $10\;\text{kW/m}^{2}/\text{K}$ in tile regions with very thick surface layers to $160\;\text{kW}/\text{m}^{2}/\text{K}$ in clean tiles. $\alpha_{top}$ is estimated as the maximum value which avoids negative heat fluxes.
4.3 Power and parallel heat flux
The heat flux perpendicular to the tile $q_{\perp}$ is obtained with THEODOR. It can be related to the plasma heat flux parallel to the magnetic field through the equation
$q_{\perp}=q_{B,\parallel}\sin\alpha+q_{B,\perp}\cos\alpha+q_{bg}$ (4.11)
where $\alpha$ is the grazing angle between the plasma and the tile (not to be confused with $\alpha_{top}$, the surface layer heat transmission factor), $q_{B,||}$ is the plasma heat flux parallel to the magnetic field, $q_{B,\perp}$ is the plasma heat flux perpendicular to the magnetic field and $q_{bg}$ is the background heat flux due to radiation. Assuming that for diverted plasmas $q_{B,||}\sin\alpha\gg q_{B,\perp}\cos\alpha$ near the strike points, the heat flux parallel to $B$ can be estimated as
$q_{\perp}=q_{\parallel}\sin\alpha+q_{bg}\implies q_{||}=\left(q_{\perp}-q_{bg}\right)/\sin\alpha$ (4.12)
where for simplicity $q_{||}\equiv q_{B,||}$.
Some tiles in TCV are not toroidally symmetric (e.g. all inner wall tiles seen by HIR, the VIR valley tile, and the TIR tiles). To evaluate the perpendicular heat flux on a toroidally symmetric tile, a correction in the grazing angle is made, yielding
$q_{\perp,\text{sym}}=q_{||}\sin\alpha_{\text{sym}}+q_{bg}$ (4.13)
If the tile is toroidally symmetric (i.e. if $\alpha_{\text{sym}}=\alpha$, as for VIR’s flat tiles), then $q_{\perp,\text{sym}}=q_{\perp}$.
The power deposited on TCV’s walls is obtained by integrating $q_{\perp,\text{sym}}$,
$P=\int_{s_{min}}^{s_{max}}\int_{0}^{2\pi}q_{\perp,\text{sym}}Rd\varphi ds=2\pi\int_{s_{min}}^{s_{max}}q_{\perp,\text{sym}}Rds$ (4.14)
where $s=z$ for the inner wall, $s=R$ for the floor, and $s=\sqrt{R^{2}+z^{2}}$ for the tilted tiles.
It is also possible to project the parallel heat flux upstream, by accounting for the total flux expansion from the upstream to the target,
$q_{||,u}=f_{x,tot}q_{||}\approx\frac{B_{u}}{B_{t}}q_{||}$ (4.15)
where $B_{u}$ is the magnetic field upstream (on the outer midplane) and $B_{t}$ is the magnetic field at the target. It is common to show the $q_{||,u}$ profile as a function of the upstream coordinate $R-R_{sep}$ where $R$ here is the major radius position upstream and $R_{sep}$ is the major radius of the separatrix.
5 Windows
TCV’s IR Equus cameras request that the windows which separate the tokamak to the camera have a high transmissivity in the range $3.7\;\mu\text{m}$ to $4.8\;\mu\text{m}$ (Table 2.1), which can be obtained for example with zinc selenide (ZnSl), germanium (Ge) or sapphire ($\text{Al}_{2}\text{O}_{3}$). Currently all TCV IR windows are made of sapphire with an anti-reflection coating. Previously used germanium windows cracked (HIR lower port, Table 2.3) or had their transmission to decay by a factor 3 with time (VIR, Table 2.3). Zinc selenide are more fragile and are therefore avoided.
6 Tiles
The tiles, together with the infrared cameras, compose the core of the IR system. Heat fluxes are inferred from the tile surface temperatures (Sec. 4.1). TCV’s tiles are made of polycrystalline graphite, which has an emissivity in the range of 0.7–0.9 (and therefore a reflectivity in the range 0.1-0.3).
6.1 Tile mass density, specific heat, diffusivity and conductivity
Using IR to infer plasma heat fluxes impinging on the tiles requires knowledge about the tile conductivity and diffusivity (Eq. (4.1) and (4.4)). For this reason, thermal properties of TCV tile samples were measured in 2023 and 2025 by S. Smith, P. Mildeova, Y. Mouasher, A. Awad, and L. Orkney from the National Physics Laboratory (NPL) of the United Kingdom. The results are shown in Fig. 6.1 and in the Tables A.2, and A.2 from the Appendix A.
The samples’ mass densities were measured at room temperature with a calibrated Mettler AE240 five-figure electronic balance, obtaining $\rho=1.836\;\text{g}/\text{cm}^{3}$ for the 2023 sample and $\rho=1.855\;\text{g}/\text{cm}^{3}$ for the 2025 one. Thermal expansion measurements (not shown) were then performed with a Linseis twin push-rod alumina dilatometer up to a maximum temperature of 1020 ºC. The thermal expansion measurements were used to evaluate the mass density at different temperatures. The mass density variation with temperature was negligible: it decreased by less than 1% in the range of 20 ºC to 500 ºC (Fig 6.1(a)).
The specific heat capacity was obtained with a TA Instruments twin-pan differential scanning calorimeter (DSC). The test-piece mass was determined with a calibrated Sartorius Secura 225D-1S electronic balance to a resolution of $0.00001\;\text{g}$. Three heating runs were performed on each sample, from 0 ºC to 430 ºC. It is seen in Fig. 6.1(b) that the specific heat increases with temperature, going from $0.7\;\text{J/g/K}$ at 20 ºC to $1.6\;\text{J/g/K}$ at $500\;\text{ºC}$ (the results above $430\;\text{ºC}$ was extrapolated).
The thermal diffusivity of the samples was determined via a laser flash analysis (LFA) using the instrument Netzsch LFA 427. The tile diffusivity decreases with the functional dependency expected by THEODOR (Eq. (4.7)), ranging from around $80\;\text{mm}^{2}/\text{s}$ at $T=0\;\text{º}\text{C}$ to $25\;\text{mm}^{2}/\text{s}$ at $500\;\text{º}\text{C}$. Fitting the data from the three samples yielded (black curve in Fig. 6.1(c))
$D_{fit}(T)=\left[10.6(2.7)\;+72.9(2.9)\left(1+\frac{T}{392(56)\;\text{ºC}}\right)^{-2}\right]\text{mm}^{2}/\text{s}$ (6.1)
where the values in parentheses ((2.7), (2.9), (56)) are the uncertainties in the parameters. Using the standard deviation between the 2023 and 2025 samples as uncertainty yielded a chi-squared of $\chi^{2}=7$, in good agreement with the number of degrees of freedom, $\text{DF}=12$. Before the NPL measurements, the old estimate used in TCV for the tile diffusivity was (yellow curve in Fig. 6.1(c))
$D_{old}(T)=\left[15\;+60\left(1+\frac{T}{53\;\text{ºC}}\right)^{-2}\right]\text{mm}^{2}/\text{s}$ (6.2)
Finally, the sample conductivity data was obtained by multiplying the mass density, specific heat, and diffusivity $k=\rho c_{p}D$ (as in Eq. (4.5)). A least-squares fit yielded (black curve in Fig. 6.1(d))
$k_{fit}(T)=104.2(1.7)\left(1+\frac{T}{2.82(0.19)\times 10^{3}\;\text{ºC}}\right)^{-2}\frac{\text{W}}{\text{m}\;\text{K}}$ (6.3)
$a_{D}$ in Eq. (4.5) was set to zero since keeping it still yielded a value compatible to 0 but with much higher uncertainties in the parameters, which is a sign of over-parametrization. The chi-squared was $\chi^{2}=9$, again in good agreement with the degrees of freedom, $\text{DF}=13$. The tile conductivity decreases from 100 W/m/K at 0ºC to 75 W/m/K at 500 ºC. The old values of thermal conductivity (yellow curve in Fig. 6.1(d)) used for TCV’s IR were smaller than the NPL results by 25% to 40% for temperatures up to 500 ºC,
$k_{old}(T)=\left[55+25\left(1+\frac{T}{92\;\text{ºC}}\right)^{-2}\right]\frac{\text{W}}{\text{m}\;\text{K}}$ (6.4)
As of January 2026, the tile diffusivity and the conductivity given by Eq. (6.1) and (6.3) are used for IR in TCV to evaluate heat fluxes with THEODOR. The effect on the estimated heat fluxes was small — less then 15% increase in the peak heat flux when compared to the old estimated thermal properties, with the difference going to 0 far from the peak.
Figure 6.1: TCV’s graphite tile properties measured by NPL. (a) Mass density. (b) Specific heat capacity. (c) Diffusivity. (d) Conductivity. Blue and circles: data from 2023 samples; orange circles: 2025 samples. Black curves: fit using the functional dependency expected by THEODOR ($f(T)=a+b(1+T/T_{0})^{-2}),$ yielding Eq. (6.1) and (6.3). Yellow stars and curves: old estimates used in TCV for the tile diffusivity and conductivity (Eq. (6.2) and (6.4)).
6.2 VIR valley tile
A new valley tile has been commissioned for the VIR system in 2025 (Fig. 6.2). The main goal of the tile is to allow the camera operate at high acquisition frequencies. This is achieved by two means, both which allow decreasing the camera integration time and therefore increasing the acquisition frequency ($f_{acq}=1/\Delta t_{frame}$, with $\Delta t_{frame}\approx\Delta t_{int}+\Delta t_{read}$ given by Eq. (2.4)). Firstly, the tile has been embedded with a heating element, which can be used to heat up the tile just before the shot. Due to the highly non-linear dependence of the infrared radiation with the tile temperature, heating it allows significant increase in the camera signal.
Secondly, the tile has parts with an inclination of 6º to increase the grazing angle, which also leads to a higher tile temperature, again allowing lower integration times. The reasoning is summarized in the following equation:
$\Delta T\propto q_{\perp}\sqrt{\Delta t}\approx\sin(\alpha)q_{\parallel}\sqrt{\Delta t}\propto\sin\alpha\approx\alpha$ (6.5)
The variation in the tile surface temperature $\Delta T$ due to a heat flux $q_{\perp}$ perpendicular to the tile applied during an interval $\Delta t$ is proportional $q_{\perp}\sqrt{\Delta t}$ [11]. The heat flux perpendicular to the tile $q_{\perp}$ is then approximately connected to the heat flux parallel to the magnetic fields through $q_{\perp}\approx q_{\parallel}\sin\alpha$. Therefore, increasing the grazing angle $\alpha$ increases $\Delta T$ almost linearly (since $\alpha\ll 1\implies\sin\alpha\approx\alpha$).
Tilting the tile in the toroidal direction by an angle $\alpha_{ti}$ then increases the temperature difference by
$\frac{\Delta T}{\Delta T_{0}}\propto\frac{\alpha_{0}+\alpha_{ti}}{\alpha_{0}}$ (6.6)
where $\alpha_{0}$ is the grazing angle between the plasma and a flat floor tile. It was chosen to incline the tile to $\alpha_{ti}=6\text{º}$, such than the increase in the temperature difference would be of a factor 3 for an initial grazing angle of $\alpha_{0}=3\text{º}$, $(3\text{º}+6\text{º})/3\text{º}=3$.
The VIR tile was designed as a valley instead of a rooftop to avoid shadowing Langmuir probes present in neighboring tiles (small dots in Fig. 6.2(b)). Additional holes with screws were added to allow the correction of vibration of the IR images even in reduced frame mode. However, it was found that the tile temperature can be modified in the vicinity of these screws, hindering the IR heat flux analysis. Therefore, if in the future a new tile is manufactured, it is recommended to not add these extra holes.
Figure 6.2: (a) VIR valley tile outside TCV. (b) TCV floor with the VIR valley tile, also showing tiles with Langmuir probes on the left. Photo (b) by Frédéric Dolizy and Luc Morier-Genoud.
6.3 TIR rooftop tiles
One pair of TIR rooftop tiles has been designed in 2025 (Fig. 6.3) to increase the camera signal by increasing the grazing angle between the plasma and the tile. The tile was designed with the following surface equation
$z_{t}(y)=\frac{y_{t}^{2}}{2R_{0}}\sin\beta+Ky_{t}\cos\beta$ (6.7)
where $R_{0}=970\;\text{mm}$ is the major radius at which the tilted tile starts, $\beta=50\text{º}$ is the angle between the tile and the floor, $y_{t}$ is the direction of the tile width ($y_{t}\in[-41.9,41.9]\;\text{mm}$ in the lowest part of the tile and $y_{t}\in[-48.8,48.8]\;\text{mm}$ at the highest), $z_{t}$ is the tile height in its referential and $K=0.0812$ is a constant calculated to transform an initial grazing angle of $\alpha=4\text{º}$ into $\alpha=7\text{º}$ by the addition of the rooftop. The derivation of Eq. (6.7) is shown in Appendix B.
One extra hole with a screw was added to the bottom of each of the tiles to allow for the correction of vibration of the IR images. However, it was found that this screw removed significant space for the TIR line of interest used for the heat flux calculation. Therefore, if in the future a new TIR rooftop tile is commissioned, it is recommended to not add this extra screw.
Figure 6.3: TIR rooftop tiles outside (a) and inside (b) TCV. Photo (b) by Frédéric Dolizy and Luc Morier-Genoud.
7 Filtering infrared plasma emission
Infrared plasma radiation (IR-PR) can distort the tile temperature measurements obtained with infrared thermography, especially at low plasma temperature and high plasma density. IR-PR can occur in three forms: bound-bound emission (atoms excited by electrons or charge exchange, emitting photons known as line emission), free-bound emission (free electron recombining with an atom and emitting the excess kinetic energy) or free-free emission (photons created during de-acceleration of free electrons, known as bremsstrahlung).
According to the NIST LIBS database, the strongest atomic line emission in the range of TCV’s IRCAM Equus cameras ($3.7\;\mu\text{m}<\lambda<4.8\;\mu\text{m}$) is the deuterium line $D_{5\rightarrow 4}$, with a wavelength of $\lambda=4051\;\text{nm}.$ To avoid measuring this line emission, two 4095 nm long wave pass filters were installed for the VIR and TIR systems (transmissivity in Fig. 7.1), whereas HIR already had a 4090 nm filter. Applying this filter also reduces continuous IR-PR, which is more energetic for lower wavelengths.
Fig. 7.2 shows the TIR view before and after mounting the filter for two identically programmed TCV standard shots, which are long-legged, ohmic, low temperature discharges (Thomson scattering pedestal electron temperature and density in the range $T_{e,ped}=10\text{\textendash}20\;\text{eV}$ and $n_{e,ped}=1\text{\textendash}2\times 10^{19}\;\text{m}^{3}$ at $t=0.7\;\text{s}$). A significant reduction in the IR plasma emission is seen with the filter. Nevertheless, the plasma leg still can be seen in the filtered image (Fig. 7.2(b)). After removing $\lambda<4095\;\text{nm}$, the strongest atomic line emission is $D_{7\rightarrow 5}$, at $\lambda=4654\;\text{nm}$ (carbon line emissions are negligible in the range $3.7\;\mu\text{m}<\lambda<4.8\;\mu\text{m}$ according to the NIST LIBS database). However, the $D_{7\rightarrow 5}$ intensity is 4 times weaker than $D_{5\rightarrow 4}$. This, together with SOLPS-ITER simulations (Sec. 4.5 of [6]), suggest that, after removing $\lambda<4095\;\text{nm}$, free-bound and free-bound emissions should dominate the measured IR-PR for TCV’s IRCAM Equus cameras.
Figure 7.1: Transmittance of the SPECTROGON LP-4095nm filter used for the VIR and TIR cameras (and similar one for HIR).
Figure 7.2: (a) Magnetic reconstruction of the TCV standard shots 77114 (blue) and 79993 (orange) at 0.7 s, showing the TIR field of view (red). (b) Unfiltered TIR view. (c) TIR view after putting the 4095 nm filter.
8 Comparing VIR and TIR heat fluxes
Examples of perpendicular heat flux profiles measured by the VIR and TIR systems are shown in Fig. 8.1 for the outer strike point of discharge 87222 (IR thermogram in Fig. 2.2(c) and Fig. 2.2(b)). Due to the increased grazing angle $\alpha$ of the TIR rooftop tile (Sec. 6.3), the TIR $q_{\perp}$ is considerably higher than its $q_{\perp,sym}$ (Fig. 8.1(a) and (c)). Nevertheless, projecting the heat flux for a toroidally symmetric tile using Eq. (4.13) gives VIR and TIR heat fluxes that agree well at low plasma density.
The target peak heat flux (Fig. 8.1(c)) decreases with time as the discharge 87222 had a density ramp, which evolves the plasma towards a detached state. Nevertheless, it is seen that the target power evaluated with Eq. (4.14) disagrees for VIR and TIR before the strike point arrives at the tile ($t<0.5\;\text{s}$) and after $t>1.0\;\text{s}$, when the VIR power increases while the TIR power decreases. This is caused by infrared plasma radiation measured by the IR cameras, which affects especially the VIR (as discussed in Sec. 7 and in further detail in Sec. 4.5 of [6]). As the density increases (Thomson scattering separatrix density $n_{e,sep}=2\times 10^{19}/\text{m}^{3}$ at $t=1.00\;\text{s}$ and $n_{e,sep}=4\times 10^{19}/\text{m}^{3}$ at $t=1.50\;\text{s}$ for 87222), the target plasma temperature lowers, increasing the plasma infrared emission, which is perceived by the VIR camera as in increase in the tile temperature. This causes the VIR estimated background heat flux and power to be overestimated (Fig. 8.1(b) and (d)).
Figure 8.1: (a) Perpendicular heat flux profiles at $t=1.00\;\text{s}$ for VIR and TIR in the discharge 87222 (same as Fig. 2.2 and Fig. 2.2). Solid lines: perpendicular heat flux projected in a toroidally symmetric tile, using Eq. (4.13). Dashed line: $q_{\perp}$ obtained with THEODOR, higher than $q_{\perp,sym}$ because of the increased grazing angle of the TIR rooftop tile (Sec. 6.3). (b) $q_{\perp,sym}$ at $t=1.50\;\text{s}$. (c) Peak of the heat flux profiles as a function of time. (d) Target power evaluated with Eq. (4.14).
9 Summary
TCV’s infrared thermography system is composed by three IRCAM Equus 81k M cameras, HIR, VIR, and TIR, whose main views map the inner wall, floor and tilted tiles of the tokamak. The camera raw data is calibrated to temperature using heated tiles and thermocouples. The temperature measurements are translated into target heat flux with the THEODOR code. New VIR valley and TIR rooftop tiles have been commissioned for fast transient studies.
The mass density, specific heat, diffusivity, and conductivity of TCV’s graphite tiles have been measured by NPL-UK in 2023 and 2025. The mass density, identified between $1.823$ and $1.855\;\text{g}/\text{cm}^{3}$ can be considered constant, decreasing by less then 1% from 20 ºC to 500 ºC. The specific heat increased with temperature, going from going from 0.7 J/g/K at 20 ºC to 1.6 J/g/K at 500 ºC. The diffusivity and conductivity agree with the functional dependence expected by THEODOR, yielding $D_{fit}(T)=\left\{10.6(2.7)\;+72.9(2.9)/\left[1+T/(392(56)\;\text{ºC})\right]^{2}\right\}\text{mm}^{2}/\text{s}$ and $k_{fit}(T)=104.2(1.7)/\left[1+T/\left(2.82(0.19)\times 10^{3}\;\text{ºC}\right)\right]\frac{\text{W}}{\text{m}\;\text{K}}.$
The infrared plasma radiation measured by the Equus cameras has been reduced with long wave pass filters, mainly to avoid the deuterium line $5\rightarrow 4$ at $4051\;\text{nm}$. However, at high plasma density the infrared plasma emission still distorts the IR camera signals in TCV, specially for the VIR system. This, together with the determination of the surface layer heat transmission factor $\alpha_{top}$, remain as the main sources of uncertainty for IR systems in TCV. Externally heating a tile should decrease the effect of parasitic infrared radiation, which is subject of future work.
10 Acknowledgements
This work has been carried out within the framework of the EUROfusion Consortium, partially funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion). The Swiss contribution to this work has been funded in part by the Swiss State Secretariat for Education, Research and Innovation (SERI). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission or SERI. Neither the European Union nor the European Commission nor SERI can be held responsible for them.
Appendix A Appendix: Tables with the NPL tile data
Tables A.2 and A.2 contain the thermal properties of TCV’s graphite tiles measured by NPL-UK, used in Fig. 6.1.
Table A.1: Mass density, specific heat, diffusivity, and conductivity for the graphite 2023 samples discussed in Sec. 6.1.
Temp [ºC] $\rho\;[\text{g/cm}^{3}]$ $c_{p}\;\text{[J/g/K]}$ $D\;[\text{mm}^{2}/\text{s}]$ $k\;[\text{J/m/K}]$
18 1.836 0.696 82.97 106
99 1.834 0.916 61.14 103
299 1.829 1.350 35.31 87.2
500 1.823 1.616 25.67 75.6
400 1.827 1.500 29.07 79.7
202 1.832 1.141 43.14 90.2
17 1.836 0.696 81.00 104
Table A.2: Mass density, specific heat, diffusivity, and conductivity for the graphite 2025 samples.
Temp [ºC] $\rho\;[\text{g/cm}^{3}]$ $c_{p}\;\text{[J/g/K]}$ $D\;[\text{mm}^{2}/\text{s}]$ $k\;[\text{J/m/K}]$
15 1.855 0.707 74.52 97.7
20 1.855 0.722 72.94 97.7
100 1.853 0.956 54.56 96.7
300 1.848 1.398 32.35 83.6
500 1.842 1.658 23.84 72.8
399 1.845 1.557 27.91 80.2
200 1.851 1.203 41.81 93.1
20 1.855 0.722 73.17 98.0
Appendix B Appendix: Derivation of the TIR rooftop tile surface
B.1 The rooftop surface equation in the tokamak frame
The surface of the simple (non-rooftop) tilted tiles satisfies
$z(x,y)=\left(x-R_{0}\right)\tan\beta$ (B.1)
where $z$ is the height of the tile, $R_{0}=970\;\text{mm}$ is the major radius at which the tilted tiles start and $\beta=50\text{º}$ is the tilted tiles angle with the floor. The reference frames used in derivation and the dimensions of the tile are described in Fig. B.1.
A toroidally symmetric tile would follow
$z=(R-R_{0})\tan\beta$ (B.2)
where $R$ is the major radius. Adding a component proportional to $R\varphi$ forms a rooftop and increases the grazing angle,
$z=(R-R_{0})\tan\beta+KR\varphi$ (B.3)
where $\varphi$ is the toroidal angle (set to 0 at the center of the tile) and $K$ is a constant set to increase a grazing angle of $\alpha_{0}=4\text{º}$ (without the rooftop) to $\alpha=7\text{º}$.
Figure B.1: (a) The three reference frames ($(x,y,z)$, $(x^{\prime},y^{\prime},z^{\prime})$, and $(x_{t},y_{t},z_{t})$) used to calculate the TIR rooftop tile equation. Note that $z=z^{\prime}$, $y=y^{\prime}=y_{t}$, and that $\beta=50\text{º}$ for the tilted tiles. (b) Dimensions of the tilted tile [$m$].
B.2 Finding the rooftop constant K
To find $K$ in Eq. (B.3), it is necessary to evaluate the grazing angle (and hence the vector normal to the surface). Let
$f(R,\varphi,z)=(R-R_{0})\tan\beta+KR\varphi-z=0$ (B.4)
Then vector normal to the surface is
$\vec{n}$ $=\nabla f(R,\varphi,z)=\frac{\partial f}{\partial R}\hat{R}+\frac{\partial f}{R\partial\varphi}\hat{\varphi}+\frac{\partial f}{\partial z}\hat{z}$
$=(\tan\beta+K\varphi)\hat{R}+K\hat{\varphi}-\hat{z}$
and the unit normal vector is
$\hat{n}=\frac{\vec{n}}{\left|\vec{n}\right|}=\frac{(\tan\beta+K\varphi)\hat{R}+K\hat{\varphi}-\hat{z}}{\left[\left(\tan\beta+K\varphi\right)^{2}+K^{2}+1\right]^{1/2}}$ (B.5)
Assuming $K\varphi\ll\tan\beta$ and $K^{2}\ll 1$,
[…]
The sine of the grazing angle between the plasma and the tile is
$\sin\alpha=\frac{\vec{B}}{B}\cdot\hat{n}$ (B.7)
where the magnetic field is
$\vec{B}=B_{R}\hat{R}+B_{\varphi}\hat{\varphi}+B_{z}\hat{z}$ (B.8)
So
$\sin\alpha$ $=\frac{1}{B}\cos\beta\left(B_{R}\tan\beta+KB_{\varphi}-B_{z}\right)$
$=\sin\alpha_{0}+\frac{B_{\varphi}}{B}K\cos\beta$
where $\sin\alpha_{0}=\frac{1}{B}\cos\beta(B_{R}\tan\beta-B_{z})$ is the grazing angle sine for $K=0$, i.e. without the rooftop. Hence
$K=\frac{B}{B_{\varphi}}\frac{\sin\alpha-\sin\alpha_{0}}{\cos\beta}$ (B.9)
Assuming $B_{\varphi}=0.998B$, $\alpha=7\text{º}$, $\alpha=4\text{º}$ and $\beta=50\text{º}$ gives
$K=0.0812$ (B.10)
Note that since $B/B_{\varphi}\approx 1$ and $\alpha\ll 1$, a floor tile (i.e. $\beta=0$) would have $K\approx\alpha-\alpha_{0}$.
B.3 TIR rooftop tile surface in the tile frame
The maximum toroidal angle variation of a tilted tile in TCV is $\varphi_{max}=1.24\text{º}\cdot\pi/180\text{º}=0.02164$. Since $\varphi_{max}\ll 1$, then $\cos\varphi_{max}=0.9998\approx 1$ and $\sin\varphi_{max}=0.02164\ll 1$. To second order in $\varphi$ and $y/x$,
$y=R\sin\varphi\approx R\varphi$ (B.11)
$R=\sqrt{x^{2}+y^{2}}=x\sqrt{1+y^{2}/x^{2}}\approx x\left(1+\frac{y^{2}}{2x^{2}}\right)=x+\frac{y^{2}}{2x}$ (B.12)
where $y_{max}/x_{0}=21/970=0.0216$ and thus $\left(y_{max}/x_{0}\right)^{2}=4.7\times 10^{-4}$. Therefore a second order approximation to the tile surface is
$z=\left(R-R_{0}\right)\tan\beta+KR\varphi\implies$ (B.13)
$\boxed{z\approx\left(x+\frac{y^{2}}{2x}-R_{0}\right)\tan\beta+Ky}$ (B.14)
It can be seen from a drawing that $z=z^{\prime}=x_{t}\sin\beta+z_{t}\cos\beta$ and $z_{t}=-x^{\prime}\sin\beta+z\cos\beta=-(x-R_{0})\sin\beta+z\cos\beta$ so
$z\cos\beta=z_{t}+(x-R_{0})\sin\beta$ (B.15)
and thus
$z\approx\left(x+\frac{y^{2}}{2x}-R_{0}\right)\tan\beta+Ky\implies$ (B.16)
$z\cos\beta=z_{t}+(x-R_{0})\sin\beta\approx\left(x+\frac{y^{2}}{2x}-R_{0}\right)\sin\beta+Ky\cos\beta\implies$ (B.17)
$z_{t}=\frac{y^{2}}{2x}\sin\beta+Ky\cos\beta$ (B.18)
Using $x-R_{0}=x^{\prime}=x_{t}\cos\beta-z_{t}\sin\beta$,
$z_{t}=\frac{1}{2}\frac{y^{2}}{R_{0}+x_{t}\cos\beta-z_{t}\sin\beta}\sin\beta+Ky\cos\beta$ (B.19)
Since $z_{t}\sin\beta\ll R_{0}$,
$z_{t}\approx\frac{1}{2}\frac{y^{2}}{R_{0}+x_{t}\cos\beta}\sin\beta+Ky\cos\beta$ (B.20)
To manufacture the tile it is easier if $z_{t}$ only depends on one coordinate. For this reason it is useful to approximate $y^{2}/(R_{0}+x_{t}\cos\beta)\approx y^{2}/R_{0}$. This gives the expression used to manufacture the tile,
$\boxed{z_{t}(y)\approx\frac{y^{2}}{2R_{0}}\sin\beta+Ky\cos\beta}$ (B.21)
A slightly more precise approximation would be $y^{2}/(R_{0}+x_{t}\cos\beta)\approx y^{2}/(R_{0}+0.5x_{t,max}\cos\beta)$.
It is possible to check if $y^{2}/(R_{0}+x_{t}\cos\beta)\approx y^{2}/R_{0}$ is reasonable. Using $y(R_{0},\varphi_{max})=41.9\;\text{mm}$, $y(R_{max},\varphi_{max})=48.8\;\text{mm}$ (Fig. B.1(b)), $x_{t}(R_{max})=247.2\;\text{mm}$, and $R_{0}\approx 970\;\text{mm}$,
$\frac{1}{2}\frac{y^{2}(R_{0},\varphi_{max})}{R_{0}}=\frac{1}{2}\frac{41.9^{2}}{970}\;\text{mm}=0.90\;\text{mm}$ (B.22)
while
$\frac{1}{2}\frac{y^{2}(R_{0},\varphi_{max})}{R_{0}+x_{t}(R_{max})\cos\beta}=\frac{1}{2}\frac{48.8^{2}}{970+247.2\cos 45\text{º}}\;\text{mm}=1.04\;\text{mm}$ (B.23)
So the maximum error is of $1.04\;\text{mm}-0.90\;\text{mm}=0.14\;\text{mm}$. Furthermore, using, $y(R_{max},\varphi_{max})=48.8\;\text{mm}$, $x_{t}(R_{max})=247.2\;\text{mm}$, and $R_{0}\approx 970\;\text{mm}$ gives
$\frac{1}{2}\frac{y^{2}(R_{max})}{R_{0}+x_{t}(R_{max})\cos\beta}\sin\beta=\frac{1}{2}\frac{48.8^{2}\sin 45\text{º}}{970+247.2\cos 45\text{º}}\;\text{mm}=0.74\;\text{mm}$ (B.24)
while the approximation is
$\frac{1}{2}\frac{y^{2}(R_{max})}{R_{0}}\sin\beta=\frac{1}{2}\frac{48.8^{2}\sin 45\text{º}}{970}\;\text{mm}=0.87\;\text{mm}$ (B.25)
Therefore the highest error committed is of $0.87\;\text{mm}-0.74\;\text{mm}=0.13\;\text{mm}$ which is close to the machining precision ($0.1\;\text{mm}$).

### 01 Apr 2026 v2

@@
 4.2 Boundary conditions and surface layer heat transmission factor
-For IR in TCV, a python version of the THEODOR code is used, which solves the differential equation (4.4) implicitly. The three boundary conditions for this THEODOR version are
-$\begin{cases}\left.\frac{\partial T}{\partial x}\right|_{\text{side}}=0\\ q_{\perp}(x,d_{tile},t)=0\\ T(x,0,t)=T_{layer}(x,t)-\frac{q_{\perp}(x,0,t-\Delta t)}{\alpha_{top}}\end{cases}$ (4.9)
-The first equation means that THEODOR assumes that the temperature derivative at the edges of the heat flux profile is zero.
-The 2nd equation assumes that the heat flux at the bottom of the tile is zero.
-Tile erosion and dust deposition can form surface layers that cause an overestimation of the actual tile temperature [8, 9, 10].
-To correct for this effect and avoid false negative heat fluxes, a surface layer heat transmission factor is introduced,
+For IR in TCV, a python version of the THEODOR code is used, which solves the differential equation (4.4) implicitly. Three boundary conditions are used for this THEODOR version. Two of them are
+$\begin{cases}\left.\frac{\partial T}{\partial x}\right|_{\text{side}}=0\\ q_{\perp}(x,d_{tile},t)=0\end{cases}$ (4.9)
+The first equation means that THEODOR presumes a vanishing temperature derivative at the edges of the heat flux profile.
+The second equation assumes that the heat flux at the bottom of the tile is zero.
+Tile erosion and particle deposition can form surface layers (SLs) that cause an overestimation of the actual tile temperature [8, 9, 10].
+In TCV, the main SL components are believed to be made of carbon (wall material), deuterium (injected gas), and boron oxides (due to boronizations).
+To account for this SL and avoid false negative heat fluxes, a surface layer heat transmission factor is introduced,
 $\alpha_{top}=\frac{k_{layer}}{d_{layer}}$ (4.10)
 where $k_{layer}$ and $d_{layer}$ are respectively the conductivity and thickness of the layer.
-The third boundary condition relates the temperature of the surface layer $T_{layer}(x,t)$ (i.e., the temperature actually measured by the IR cameras), to the tile temperature just below the surface layer, $T(x,0,t)$.
-$q_{\perp}(x,0,t-\Delta t)$ is used since $q_{\perp}(x,0,t)$ is not known at the time where the boundary condition is computed.
+Differently from previous IDL versions, the Python THEODOR does not use use the boundary condition $T(x,0,t)=T_{layer}(x,t)-\frac{q_{\perp}(x,0,t-\Delta t)}{\alpha_{top}}$ (as it was the case in Eq.
+(4.11) of [6]).
+Instead, an actual layer is modeled on top of the tile using $\alpha_{top}$ as input.
+The SL thickness is assumed to be of $d_{layer}=10\;\mu\text{m}$, although in reality this value could change depending on the tile exposition to the plasma.
+The SL conductivity is then obtained with $k_{layer}=\alpha_{top}d_{layer}$.
+The SL mass density times heat capacity, $\rho_{layer}c_{p,layer}$, is supposed to be 100 times smaller than the top of the tile just below the SL, which is then used to estimate the SL diffusivity, $D_{layer}=k_{layer}/(\rho_{layer}c_{p,layer})$.
 Values for $\alpha_{top}$ in TCV range from $10\;\text{kW/m}^{2}/\text{K}$ in tile regions with very thick surface layers to $160\;\text{kW}/\text{m}^{2}/\text{K}$ in clean tiles.
@@
 10 Acknowledgements
+The authors thank Michael Faitsch and Dirk Stieglitz for insightful discussions and continued support with THEODOR.
 This work has been carried out within the framework of the EUROfusion Consortium, partially funded by the European Union via the Euratom Research and Training Programme (Grant Agreement No 101052200 — EUROfusion).